# Interpretable and flexible non-intrusive reduced-order models using reproducing kernel Hilbert spaces

### v1

@@ 6 Numerical results
where $\mathbf{q}(t),\mathbf{q}_{0}({\boldsymbol{\mu}})\in{}^{n_{q}}$ and $\mathbf{A}\in{}^{n_{q}\times n_{q}}$. We use $n_{q}=256$ spatial degrees of freedom in this experiment. To collect training data, we sample $M=10$ initial conditions corresponding to $10$ Latin hypercube samples from the parameter domain ${\cal D}=[0.25,0.35]\times[0.05,0.15]$ and integrate the FOM eq. 6.2 using a fully implicit variable-order backwards difference formula (BDF) time stepper with quasi-constant step size, executed with scipy.interpolate.solve_ivp() in Python [60, 55]. The solution is recorded at $n_{t}=256$ equally spaced time instances after the initial condition, resulting in $M(n_{t}+1)=2570$ total training snapshots. We also solve the FOM at the testing parameter value $\bar{{\boldsymbol{\mu}}}=(0.3,0.1)$, which is not included in the training set. Figure 1 plots the FOM states for two training parameter values and the testing parameter value.
@@
where $\tilde{\mathbf{c}}\in{}^{r}$ and $\tilde{\mathbf{A}}\in{}^{r\times r}$, whereas the intrusive QM ROM has the form
@@
with $\tilde{\mathbf{c}}\in{}^{r}$, $\tilde{\mathbf{A}}\in{}^{r\times r}$, and $\tilde{\mathbf{H}}\in{}^{r\times r^{2}}$. For both POD and QM, we construct feature map Kernel ROMs and OpInf ROMs with the corresponding intrusive ROM structure. The underlying feature maps ${\boldsymbol{\phi}}$ and weighting matrices $\mathbf{G}$ are listed in Table 2. Note that the second diagonal block in the weight $\mathbf{G}$ for QM is scaled by $\left\|\mathbf{W}\right\|_{F}$ to account for the fact that $\tilde{\mathbf{H}}$ in the intrusive QM ROM eq. 6.4 also depends on $\mathbf{W}$. We also construct an RBF Kernel ROM using a Gaussian kernel-generating RBF $\psi$ (see Table 1) with fixed shape parameter $\epsilon=10^{-1}$. This ROM has the same evolution equations in the POD and QM cases, since the compression map $\mathbf{h}$ is the same in both instances, but we report results for both POD and QM decompression maps $\mathbf{g}$. Finally, we construct hybrid Kernel ROMs using the POD feature map from Table 2 with weighting coefficient $c_{\phi}=1$ and a Gaussian RBF kernel with $\epsilon=0.1$ and weighting coefficient $c_{\psi}=10^{-3}$, yielding ROMs with the following structure:
@@
[…]
@@
Here, $\nu>0$ is the viscosity, which we set to $\nu=10^{-2}$ for our experiments. Solutions to this system are characterized by sharp gradients along an advection front. Just as in the previous problem, we consider parameterized Gaussian initial conditions with center $\mu_{1}\in[0.25,0.35]$ and $\mu_{2}\in[0.05,0.15]$, set the final time to $T=1$, use $n_{q}=n_{t}=256$ spatial degrees of freedom and temporal observations, and draw $M=10$ latin hypercube samples of the parameters ${\boldsymbol{\mu}}=[~{}\mu_{1}~{}~{}\mu_{2}~{}]^{\mathsf{T}}$ to use for generating training data. The spatial discretization uses uniform centered finite differences, yielding a quadratic FOM of the form
@@
where $\mathbf{q}(t),\mathbf{q}_{0}({\boldsymbol{\mu}})\in{}^{n_{q}}$, $\mathbf{A}\in{}^{n_{q}\times n_{q}}$, and $\mathbf{H}\in{}^{n_{q}\times n_{q}^{2}}$. We again use a BDF time integrator to solve the FOM (and constructed ROMs) at the parameter samples, resulting in $M=10$ trajectories of $n_{t}+1=257$ snapshots each. The FOM states for a few parameter values are displayed in Figure 5.
POD ${\boldsymbol{\phi}}(\hat{\mathbf{q}})=\begin{bmatrix}1\\ \hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\end{bmatrix},\qquad\mathbf{G}=\frac{1}{1+r+r^{2}}\mathbf{I}_{1+r+r^{2}}$
QM ${\boldsymbol{\phi}}(\hat{\mathbf{q}})=\begin{bmatrix}1\\ \hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\end{bmatrix},\qquad\mathbf{G}=\begin{bmatrix}\mathbf{I}_{1+r+r^{2}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\left\|\mathbf{W}\right\|_{F}\mathbf{I}_{r^{3}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\left\|\mathbf{W}\right\|_{F}^{2}\mathbf{I}_{r^{4}}\\ \end{bmatrix}$
@@
For both POD and QM, we use $\bar{\mathbf{q}}=\bf 0$, hence the intrusive POD ROM takes the quadratic form eq. 3.15, whereas the intrusive QM ROM has the quartic form eq. 3.16. For this problem, we apply the kernel input normalization discussed in Remark 2.1, which is helpful for balancing the contribution of higher-order terms. We therefore construct feature map Kernel ROMs to mirror the structure of the intrusive models, with the addition of a constant term that arises due to the input scaling (see Appendix A), by using the feature maps and weighting matrices listed in Table 3. Similar to before, we learn Gaussian RBF Kernel ROMs with fixed shape parameter $\epsilon=10^{-1}$ and hybrid Kernel ROMs using the POD feature map from Table 3 with weighting coefficient $c_{\phi}=1$ and a Gaussian RBF kernel with $\epsilon=0.1$ and weighting coefficient $c_{\psi}=10^{-3}$, which result in ROMs of the form
@@
Figure 6 reports results for various reduced dimensions $r$. All POD ROM errors are nearly identical to the POD projection error. For the QM ROMs, the OpInf and Kernel ROMs have very similar performance for $r\leq 14$. The feature map Kernel ROM errors plateau for $r>14$, while the OpInf and RBF Kernel ROMs plateau for $1816<r<$ 18 and increase slightly for $r=19,20$. Notably, the hybrid Kernel ROM continues to match the projection and intrusive ROM error as $r$ increases, indicating that the RBF term in eq. 6.11 acts as a more accurate closure term for the ROM dynamics at larger values of $r$ compared to the cubic and quartic nonlinearities of the OpInf and FM Kernel ROMs. Unlike the advection-diffusion case, the QM projection and intrusive ROM errors are notably lower than the corresponding POD errors, and thus QM dimension reduction may be beneficial for this problem.
@@
We next compute the error bound from Theorem 5.1 for the FM Kernel ROMs for $r=6,12$. The quantities $\|\tilde{\mathbf{f}}+{\boldsymbol{\delta}}\|_{{\cal H}_{K}^{r}}$, $\delta(s)$, $\Lambda_{\mathbf{M}}[\mathbf{f}](\mathbf{g}(\hat{\mathbf{q}}(s)))$ are estimated in the same way as in the advection-diffusion case. We use feature map Kernel ROMs corresponding to the quadratic and quartic feature maps in Table 3 and examine the cases when the chosen feature map does not match the true projection-based ROM form, i.e. POD with a quartic feature map and QM with a quadratic feature map. Figure 7 displays the results and shows that the computed error estimate again bounds the true error without dramatically overestimating it. In the POD cases with quadratic ROMs, the $\alpha_{P}$ term dominates the error bound contribution, while the $\alpha_{K}$ term is negligible in the $r=6$ case, but less negligible in the $r=12$ case. For the QM quartic ROMs, the $\alpha_{P}$ and $\alpha_{K}$ terms contribute similarly to the error bound evaluation. The $r=6$ case contrasts with the advection-diffusion QM ROM case in that the $\alpha_{K}$ term is non-negligible despite having a model form that should reproduce the projection-based ROM model form. We again compute the error bound for a QM ROM with the cubic and quartic terms removed, which come from the quadratic part of $\mathbf{g}$, resulting in a QM quadratic ROM. As in the advection-diffusion example, the $\alpha_{K}$ term decreases significantly, which may indicate that a quadratic model form may be the better choice for a QM Burgers ROM. To again test if an incorrect model form significantly increases $\alpha_{K}$, we compute a POD quartic ROM and observe that $\alpha_{K}$ is much larger than for the POD quadratic ROM, as expected. This further evidences that a larger $\alpha_{K}$ contribution may indicate that a non-optimal model form is being used for the Kernel ROM.
@@
The initial condition is given by a classical Riemann problem as follows. The spatial domain is divided into four quadrants with a vertical dividing line at $x=0.8$ and a horizontal dividing line at $y=0.8$. The initial pressure is set to $p_{BL}=0.029$ in the bottom left quadrant; in the top right quadrant, the initial velocities are fixed at $u_{TR}=v_{TR}=0$, and the initial density is $\rho_{TR}=1.5$. We parameterize the initial condition by setting the upper-right quadrant pressure to $p_{TR}\in\left\{0.5,0.75,1.0,1.25,1.5\right\}$ and compute remaining quantities following the relations in [53, Configuration 3]. For testing, we consider the initial upper-right quadrant pressure to $\bar{p}_{TR}=1.125$. In every case, the discontinuities of the initial condition propagate through the domain, a highly challenging scenario for projection-based model reduction.
@@
Figure 9: First $2000$ normalized singular values $\sigma_{k}/\sigma_{1}$ for the 2D Euler example.
@@
This FOM is not formed explicitly, but it motivates an appropriate structure for feature map Kernel ROMs using POD or QM. In both cases, we set $\bar{\mathbf{q}}$ to the average training snapshot and apply the kernel input normalization from Remark 2.1, leading to a POD ROM structure
@@
We plot the error eq. 6.20 versus the reduced dimension $r$ for the POD OpInf, feature map Kernel, RBF Kernel, and hybrid Kernel ROMs in Figure 10. For $r=5,10,15$, each of the ROMs obtain nearly identical performance. For $r>15$, the projection error and the Kernel ROM errors plateau, with the Kernel ROMs yielding a $<2\%$ difference in error compared to the projection error. The Hybrid and FM Kernel ROMs have nearly identical errors, while the RBF yields slightly different but very similar errors. The OpInf ROM increases slightly in error for $r>15$, yet still obtains errors within a few percent of the projection error. We note that the plateauing of the ROM and projection errors for the tested ROM sizes is expected since the singular value decay is slow, as shown in Figure 9.
@@ Appendix A Quadratic systems with QM approximations
for nonzero $\bar{\mathbf{q}}\in{}^{n_{q}}$ and $\mathbf{W}\in{}^{n_{q}\times r^{2}}$. Specifically, we show that a nonzero reference vector $\bar{\mathbf{q}}$ causes a constant term to appear in the ROM dynamics.
@@
$\mathbf{g}(\tilde{\mathbf{q}})\otimes\mathbf{g}(\tilde{\mathbf{q}})$ $=\big{(}\bar{\mathbf{q}}+\mathbf{V}\tilde{\mathbf{q}}+\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}]\big{)}\otimes\big{(}\bar{\mathbf{q}}+\mathbf{V}\tilde{\mathbf{q}}+\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}]\big{)}$
$=\bar{\mathbf{q}}\otimes\bar{\mathbf{q}}+\bar{\mathbf{q}}\otimes(\mathbf{V}\tilde{\mathbf{q}})+\bar{\mathbf{q}}\otimes(\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}])$
$\qquad+(\mathbf{V}\tilde{\mathbf{q}})\otimes\bar{\mathbf{q}}+(\mathbf{V}\tilde{\mathbf{q}})\otimes(\mathbf{V}\tilde{\mathbf{q}})+(\mathbf{V}\tilde{\mathbf{q}})\otimes(\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}])$
$\qquad+(\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}])\otimes\bar{\mathbf{q}}+(\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}])\otimes(\mathbf{V}\tilde{\mathbf{q}})+(\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}])\otimes(\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}])$
@@
$\mathbf{f}(\mathbf{g}(\tilde{\mathbf{q}}))$ $=\mathbf{A}\big{(}\bar{\mathbf{q}}+\mathbf{V}\tilde{\mathbf{q}}+\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}]\big{)}+\mathbf{H}[\mathbf{g}(\tilde{\mathbf{q}})\otimes\mathbf{g}(\tilde{\mathbf{q}})]$
@@
$\frac{\textrm{d}}{\textrm{d}t}\tilde{\mathbf{q}}(t)&=\mathbf{V}^{\mathsf{T}}\mathbf{f}(\mathbf{g}(\tilde{\mathbf{q}}(t)))\\ &=\tilde{\mathbf{c}}+\tilde{\mathbf{A}}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{H}}_{2}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{3}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{4}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)],\end{aligned}$ (A.1a)
@@
[…]
@@
with $\tilde{\mathbf{c}}=\bf 0$. However, the Kernel ROM targets a shifted and scaled reduced state $\hat{\mathbf{q}}(t)={\boldsymbol{\Sigma}}^{-1}(\tilde{\mathbf{q}}(t)-\bar{\mathbf{x}})$ for some ${\boldsymbol{\Sigma}}\in{}^{r\times r}$ and $\bar{\mathbf{x}}\in{}^{r}$, which evolves according to
@@
$\hat{\mathbf{c}}$ $={\boldsymbol{\Sigma}}^{-1}\big{(}\tilde{\mathbf{A}}\bar{\mathbf{x}}+\tilde{\mathbf{H}}_{2}[\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}]+\tilde{\mathbf{H}}_{3}[\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}]+\tilde{\mathbf{H}}_{4}[\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}]\big{)}\in{}^{r},$
@@
$\qquad\qquad+\tilde{\mathbf{H}}_{4}({\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}})\big{)}\in{}^{r\times r},$
@@
$\qquad\qquad\qquad+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}+\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}})\big{)}\in{}^{r\times r^{2}},$

### v2

@@ 6 Numerical results
where $\mathbf{q}(t),\mathbf{q}_{0}({\boldsymbol{\mu}})\in\real^{n_{q}}$ and $\mathbf{A}\in\real^{n_{q}\times n_{q}}$. We use $n_{q}=256$ spatial degrees of freedom in this experiment. To collect training data, we sample $M=10$ initial conditions corresponding to $10$ Latin hypercube samples from the parameter domain ${\cal D}=[0.25,0.35]\times[0.05,0.15]$ and integrate the FOM eq. 6.2 using a fully implicit variable-order backwards difference formula (BDF) time stepper with quasi-constant step size, executed with scipy.interpolate.solve_ivp() in Python [61, 56]. The solution is recorded at $n_{t}=256$ equally spaced time instances after the initial condition, resulting in $M(n_{t}+1)=2570$ total training snapshots. We also solve the FOM at the testing parameter value $\bar{{\boldsymbol{\mu}}}=(0.3,0.1)$, which is not included in the training set. Figure 1 plots the FOM states for two training parameter values and the testing parameter value.
@@
where $\tilde{\mathbf{c}}\in\real^{r}$ and $\tilde{\mathbf{A}}\in\real^{r\times r}$, whereas the intrusive QM ROM has the form
@@
with $\tilde{\mathbf{c}}\in\real^{r}$, $\tilde{\mathbf{A}}\in\real^{r\times r}$, and $\tilde{\mathbf{H}}\in\real^{r\times r^{2}}$. For both POD and QM, we construct feature map Kernel ROMs and OpInf ROMs with the corresponding intrusive ROM structure. The underlying feature maps $\boldsymbol{\phi}$} and weighting matrices $\mathbf{G}$ are listed in Table 2. Note that the second diagonal block in the weight $\mathbf{G}$ for QM is scaled by $\left\|\mathbf{W}\right\|_{F}$ to account for the fact that $\tilde{\mathbf{H}}$ in the intrusive QM ROM eq. 6.4 also depends on $\mathbf{W}$. We also construct an RBF Kernel ROM using a Gaussian kernel-generating RBF $\psi$ (see Table 1) with fixed shape parameter $\epsilon=10^{-1}$. This ROM has the same evolution equations in the POD and QM cases, since the compression map $\mathbf{h}$ is the same in both instances, but we report results for both POD and QM decompression maps $\mathbf{g}$. Finally, we construct hybrid Kernel ROMs using the POD feature map from Table 2 with weighting coefficient $c_{\phi}=1$ and a Gaussian RBF kernel with $\epsilon=0.1$ and weighting coefficient $c_{\psi}=10^{-3}$, yielding ROMs with the following structure:
@@
$\gamma=\underset{\gamma}{\arg\min}\sum_{\ell=1}^{M}\sum_{k=0}^{n_{t}}\big\|\hat{\mathbf{q}}_{k}^{(\ell)}-\hat{\mathbf{q}}(t_{k};{\boldsymbol{\mu}}_{\ell},\gamma)\big\|_{2},$ (6.7)
@@
Here, $\nu>0$ is the viscosity, which we set to $\nu=10^{-4}$ for our experiments. Solutions to this system are characterized by sharp gradients along an advection front. Just as in the previous problem, we consider parameterized Gaussian initial conditions with center $\mu_{1}\in[0.25,0.35]$ and $\mu_{2}\in[0.05,0.15]$, set the final time to $T=1$, use $n_{q}=n_{t}=256$ spatial degrees of freedom and temporal observations, and draw $M=10$ latin hypercube samples of the parameters ${\boldsymbol{\mu}}=[~\mu_{1}~~\mu_{2}~]^{\mathsf{T}}$ to use for generating training data. The spatial discretization uses a uniform centered finite difference for the second derivative term and an upwind finite difference for the first derivative term, yielding a quadratic FOM of the form
@@
where $\mathbf{q}(t),\mathbf{q}_{0}({\boldsymbol{\mu}})\in\real^{n_{q}}$, $\mathbf{A}\in\real^{n_{q}\times n_{q}}$, and $\mathbf{H}\in\real^{n_{q}\times n_{q}^{2}}$. We again use a BDF time integrator to solve the FOM (and constructed ROMs) at the parameter samples, resulting in $M=10$ trajectories of $n_{t}+1=257$ snapshots each. The FOM states for a few parameter values are displayed in Figure 5.
POD ${\boldsymbol{\phi}}(\hat{\mathbf{q}})=\begin{bmatrix}\hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\end{bmatrix},\qquad\mathbf{G}=\frac{1}{r+r^{2}}\mathbf{I}_{r+r^{2}}$
QM ${\boldsymbol{\phi}}(\hat{\mathbf{q}})=\begin{bmatrix}\hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\\ \hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}\end{bmatrix},\qquad\mathbf{G}=\begin{bmatrix}\mathbf{I}_{r+r^{2}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\left\|\mathbf{W}\right\|_{F}\mathbf{I}_{r^{3}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\left\|\mathbf{W}\right\|_{F}^{2}\mathbf{I}_{r^{4}}\\ \end{bmatrix}$
@@
For both POD and QM, we use $\bar{\mathbf{q}}=\bf 0$, hence the intrusive POD ROM takes the quadratic form eq. 3.15, whereas the intrusive QM ROM has the quartic form eq. 3.16. We therefore construct feature map Kernel ROMs to mirror the structure of the intrusive models by using the feature maps and weighting matrices listed in Table 3. While the intrusive ROM informs the choice of feature map, we omit a comparison to the intrusive ROM since Galerkin ROMs are known to struggle in advection-dominated settings—often requiring additional stabilization or the usage of a least-squares Petrov-Galerkin projection (see, e.g., [14])—and focus instead on comparisons between the Kernel and OpInf ROMs and the projection error. Similar to before, we learn Gaussian RBF Kernel ROMs with fixed shape parameter $\epsilon=10^{-1}$ and hybrid Kernel ROMs using the POD feature map from Table 3 with weighting coefficient $c_{\phi}=1$ and a Gaussian RBF kernel with $\epsilon=0.1$ and weighting coefficient $c_{\psi}=10^{-3}$, which result in ROMs of the form
@@
Figure 6 reports results for various reduced dimensions $r$. The POD ROM errors and the POD projection error are nearly identical for $r\leq 14$. The OpInf POD ROM plateaus in error for $r\geq 15$, while the FM and hybrid Kernel ROMs plateau for $r\geq 16$ with slightly improved performance compared to OpInf. The RBF Kernel ROM errors are nearly identical to the projection errors for each value of $r$. In the QM case, the ROMs yield similar errors to the QM projection error for $r\leq 6$. For $r>6$, each of the QM ROMs have similar performance, and roughly plateau in error with nearly an order of magnitude difference in comparison to the QM projection error. The QM OpInf and FM Kernel ROM errors slightly decrease monotonically as $r$ increases, except in the range $8\leq r\leq 12$. The Hybrid Kernel ROM has similar performance to the FM Kernel and OpInf ROMs, but has slightly worse error for $r\geq 17$. The RBF Kernel ROM errors have non-monotonic behavior as $r$ increases, but are the lowest for each value of $r$. The QM ROMs overall obtain smaller errors than the POD projection error for $r\leq 16$, but slightly larger errors for $r>16$. These results indicate that while QM is effective at decreasing the projection error relative to POD, its benefits are limited when integrated into OpInf or Kernel ROMs in this particular experiment.
@@
We next compute the error bound from Theorem 5.1 for the FM Kernel ROMs for $r=6,12$. The quantities $\|\tilde{\mathbf{f}}+{\boldsymbol{\delta}}\|_{{\cal H}_{K}^{r}}$, $\delta(s)$, $\Lambda_{\mathbf{M}}[\mathbf{f}](\mathbf{g}(\hat{\mathbf{q}}(s)))$ are estimated in the same way as in the advection-diffusion case. We use feature map Kernel ROMs corresponding to the quadratic and quartic feature maps in Table 3 and examine the cases when the chosen feature map does not match the true projection-based ROM form, i.e. POD with a quartic feature map and QM with a quadratic feature map. Figure 7 displays the results and shows that the computed error estimate again bounds the true error, but is much larger than in the advection-diffusion case. We hypothesize that this difference in magnitude is related to the viscosity parameter $\nu=10^{-4}$. In the POD cases with quadratic ROMs, the $\alpha_{P}$ term dominates the error bound contribution, while the $\alpha_{K}$ term is negligible in the $r=6$ case, but less negligible in the $r=12$ case. For the QM quartic ROMs, in the $r=6$ case, the $\alpha_{P}$ term dominates the error bound evaluation, while for $r=12$, the $\alpha_{P}$ and $\alpha_{K}$ terms contribute similarly. We again compute the error bound for a QM ROM with the cubic and quartic terms removed, which come from the quadratic part of $\mathbf{g}$, resulting in a QM quadratic ROM. As in the advection-diffusion example, the $\alpha_{K}$ term decreases significantly, which may indicate that a quadratic model form may be the better choice for a QM Burgers ROM. To again test if an incorrect model form significantly increases $\alpha_{K}$, we compute a POD quartic ROM and observe that $\alpha_{K}$ is much larger than for the POD quadratic ROM, as expected. This further evidences that a larger $\alpha_{K}$ contribution may indicate that a non-optimal model form is being used for the Kernel ROM.
@@
The initial condition is given by a classical Riemann problem as follows. The spatial domain is divided into four quadrants with a vertical dividing line at $x=0.8$ and a horizontal dividing line at $y=0.8$. The initial pressure is set to $p_{BL}=0.029$ in the bottom left quadrant; in the top right quadrant, the initial velocities are fixed at $u_{TR}=v_{TR}=0$, and the initial density is $\rho_{TR}=1.5$. We parameterize the initial condition by setting the upper-right quadrant pressure to $p_{TR}\in\left\{0.5,0.75,1.0,1.25,1.5\right\}$ and compute remaining quantities following the relations in [54, Configuration 3]. For testing, we consider the initial upper-right quadrant pressure to $\bar{p}_{TR}=1.125$. In every case, the discontinuities of the initial condition propagate through the domain, a highly challenging scenario for projection-based model reduction.
@@
Figure 9: First $2000$ normalized singular values $\sigma_{k}/\sigma_{1}$ for the 2D Euler example. The inset figure shows the first $60$ normalized singular values to show where the singular value decay begins to slow.
@@
For further details, see, e.g., [44]. This FOM is not formed explicitly, but it motivates an appropriate structure for feature map Kernel ROMs using POD or QM. In both cases, we set $\bar{\mathbf{q}}$ to the average training snapshot and apply the kernel input normalization from Remark 2.1, leading to a POD ROM structure
@@
We plot the error eq. 6.20 versus the reduced dimension $r$ for the POD OpInf, feature map Kernel, RBF Kernel, and hybrid Kernel ROMs in Figure 10. For $r=5,10,15$, each of the ROMs obtain nearly identical performance. For $r>15$, the projection error and the Kernel ROM errors plateau, with the Kernel ROMs yielding a $2\%$ difference in error compared to the projection error. The Hybrid and FM Kernel ROMs have nearly identical errors, while the RBF yields slightly different but very similar errors. The OpInf ROM increases slightly in error for $r>15$, yet still obtains errors within a few percent of the projection error. We note that the plateauing of the ROM and projection errors for the tested ROM sizes is expected since the singular value decay is slow, as shown in Figure 9.
@@ Appendix A Quadratic systems with QM approximations
for nonzero $\bar{\mathbf{q}}\in\real^{n_{q}}$ and $\mathbf{W}\in\real^{n_{q}\times r^{2}}$. Specifically, we show that a nonzero reference vector $\bar{\mathbf{q}}$ causes a constant term to appear in the ROM dynamics.
@@
[…]
@@
[…]
@@
$\begin{aligned} \frac{\textrm{d}}{\textrm{d}t}\tilde{\mathbf{q}}(t)&=\mathbf{V}^{\mathsf{T}}\mathbf{f}(\mathbf{g}(\tilde{\mathbf{q}}(t)))\\ &=\tilde{\mathbf{c}}+\tilde{\mathbf{A}}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{H}}_{2}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{3}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{4}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)],\end{aligned}$ (A.1a)
@@
$\begin{aligned} \tilde{\mathbf{c}}&=\mathbf{V}^{\mathsf{T}}(\mathbf{A}\bar{\mathbf{q}}+\mathbf{H}[\bar{\mathbf{q}}\otimes\bar{\mathbf{q}}])\in\real^{r},\\ \tilde{\mathbf{A}}&=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}+\mathbf{V}^{\mathsf{T}}\mathbf{H}(\bar{\mathbf{q}}\otimes\mathbf{V}+\mathbf{V}\otimes\bar{\mathbf{q}})\in\real^{r\times r},\\ \tilde{\mathbf{H}}_{2}&=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{W}+\mathbf{V}^{\mathsf{T}}\mathbf{H}(\bar{\mathbf{q}}\otimes\mathbf{W}+\mathbf{V}\otimes\mathbf{V}+\mathbf{W}\otimes\bar{\mathbf{q}})\in\real^{r\times r^{2}},\\ \tilde{\mathbf{H}}_{3}&=\mathbf{V}^{\mathsf{T}}\mathbf{H}(\mathbf{V}\otimes\mathbf{W}+\mathbf{W}\otimes\mathbf{V})\in\real^{r\times r^{3}},\\ \tilde{\mathbf{H}}_{4}&=\mathbf{V}^{\mathsf{T}}\mathbf{H}(\mathbf{W}\otimes\mathbf{W})\in\real^{r\times r^{4}}.\end{aligned}$ (A.1b)
@@
with $\tilde{\mathbf{c}}=\bf 0$. However, the Kernel ROM targets a shifted and scaled reduced state $\hat{\mathbf{q}}(t)={\boldsymbol{\Sigma}}^{-1}(\tilde{\mathbf{q}}(t)-\bar{\mathbf{x}})$ for some ${\boldsymbol{\Sigma}}\in\real^{r\times r}$ and $\bar{\mathbf{x}}\in\real^{r}$, which evolves according to
@@
$\hat{\mathbf{c}}$ $={\boldsymbol{\Sigma}}^{-1}\big(\tilde{\mathbf{A}}\bar{\mathbf{x}}+\tilde{\mathbf{H}}_{2}[\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}]+\tilde{\mathbf{H}}_{3}[\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}]+\tilde{\mathbf{H}}_{4}[\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}]\big)\in\real^{r},$
@@
$\qquad\qquad+\tilde{\mathbf{H}}_{4}({\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}})\big)\in\real^{r\times r},$
@@
$\qquad\qquad\qquad+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}+\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}+\bar{\mathbf{x}}\otimes\bar{\mathbf{x}}\otimes{\boldsymbol{\Sigma}}\otimes{\boldsymbol{\Sigma}})\big)\in\real^{r\times r^{2}},$